# A Non-Relativistic Limit for Heterotic Supergravity and its Gauge Lagrangian

### v3

@@ 5.1 Outlook
As promising avenue for future work, it would be very interesting to explore the incorporation of $\alpha´$-corrections to both the bosonic and the heterotic supergravity  [^2] See [32] for a pedagogical review on this topic and references therein. in the NR limit. Since the Riemann tensor contains divergences, one should study whether these are controlled by the $\hat{H}_{\mu\nu\rho}$ (or $\bar{H}_{\mu\nu\rho}$ in the heterotic case) contributions, as happens at the two derivative level. The covariant way of writing the Ricci scalar using the covariant derivative of $\tau_{\mu}{}^{a}$ and $\tau^{\mu}{}_{a}$ could be easily extended in order to incorporate the spin connection $\omega_{\mu ab}$, so that the action is also invariant under the $SO(1,1)$ transformations. In this way, the higher-derivative contributions could be written in a more compact form, since we expect large contributions if the curvatures are not chosen in an optimal way.
While DFT has been proved a powerful tool to read the expansion of the A-field [22], taking the NR limit at the DFT level including higher-derivative terms is probably the best strategy to read these corrections and then parametrize them to the supergravity level.
@@ Acknowledgements
E.L. is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
Appendix A Details on the cancellation of divergences: gauge case
The field strength $F_{\mu\nu}{}^{i}$, after expanding $A_{\mu}{}^{i}$ considering (1.12) is given by
$F^{(1)}_{\mu\nu}{}^{i}=\alpha_{-}^{i}\partial_{\mu}{\tau_{\nu}^{-}}+\partial_{\mu}{\alpha_{-}^{i}}\tau_{\nu}^{-}-\alpha_{-}^{i}\partial_{\nu}{\tau_{\mu}^{-}}-\partial_{\nu}{\alpha_{-}^{i}}\tau_{\mu}^{-}\,,$ (A.1)
$F^{(0)}_{\mu\nu}{}^{i}=+\alpha_{-}^{j}\eta^{ik}\tau_{\mu}^{-}a_{\nu}^{i1}f_{jki1}-\alpha_{-}^{j}\eta^{ik}\tau_{\nu}^{-}a_{\mu}^{i1}f_{jki1}\,,$ (A.2)
$F^{(-1)}_{\mu\nu}{}^{i}=\partial_{\mu}{a_{\nu}^{i}}-\partial_{\nu}{a_{\mu}^{i}}\,,$ (A.3)
$F^{(-2)}_{\mu\nu}{}^{i}=-\eta^{ij}a_{\mu}^{k}a_{\nu}^{i1}f_{jki1}\,,$ (A.4)
where the notation $F^{(i)}_{\mu\nu}{}^{i}$ denotes the terms of order $c^{i}$ in the expansion. The contraction of this field strength with two inverse metrics, $F^{\mu\nu i}=F_{\rho\sigma}g^{\mu\rho}g^{\nu\sigma}$ yields the expression,
$F^{\mu\nu i}=F^{(1)\mu\nu i}+\mathcal{O}\big{(}\frac{1}{c}\big{)}\,,$ (A.5)
@@ Appendix A Details on the cancellation of divergences: gauge case
$F^{(1)\mu\nu i}=\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\sigma}h^{\nu\rho}\,.$ (A.6)
As the reader can easily check, the only contraction which produces non-vanishing divergences is given by $F_{(1)\mu\nu}{}^{i}F^{(1)\mu\nu j}\eta_{ij}$, which leads to the result (2.5).
For the $-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}$ contributions, an important observation is that
@@
while the same object when all the indices raised by the inverse metric has the following form,

### v5

@@ 5.1 Outlook
This paper opens several promising directions for future work. First, it would be very interesting to study what happens to the equations of motion under the NR limit, especially from the DFT point of view. Even in the bosonic case, these equations are finite by construction (since the generalized metric is finite) and therefore there are no equations missing in the double geometry. However, using the strong constraint to break the duality group could be related to the analysis of the equations of motion performed in [13]. When the analysis is performed directly from the NR supergravity perspective, the Poisson equation is missed and therefore it is not very clear how this equation is lost upon parametrization of the generalized metric. Once this mechanism is understood for the bosonic case, the inclusion of the gauge group should be straightforward.
Second, a very promising line of research is behind the cancellation of the Green-Schwarz mechanism under the NR expansion shown in this paper. Using the field redefinition given in (3.17), we have shown that the Green-Schwarz mechanism for gauge transformations can be trivialized. The key point here is the presence of the vector $\alpha_{-i}$, instead of the connection $a_{\mu i}$, which allowed us to remove this transformation. This is a strong indication that the anomaly cancellation in the heterotic string could not need a fixed gauge group to work under the NR limit. If this is the case, the present results open the possibility of new families of heterotic strings, considering arbitrary gauge groups. A detailed analysis of the anomaly cancellation is therefore an important step towards the construction of NR heterotic string theory.
Third, another potential continuation is to include ${\cal N}=1$ supersymmetry in this framework, extending the results in [15]. The inclusion of this symmetry in heterotic DFT requires the use of the generalized vielbein and, therefore, it is not straightforward to construct a finite action in the double geometry. However, once this construction is understood, the inclusion of the gaugino dynamics will be obtained by promoting the duality group from $O(D,D)$ to $O(D,D+N)$. So, this is a worthy program to explore NR heterotic supergravity in the presence of fermions.
Finally, a very ambitious continuation would be the incorporation of $\alpha$´-corrections to both the bosonic and the heterotic supergravity  [^2] See [32] for a pedagogical review on this topic and references therein. in the NR limit. Since the Riemann tensor contains divergences, one should study whether these are controlled by the $\hat{H}_{\mu\nu\rho}$ (or $\bar{H}_{\mu\nu\rho}$ in the heterotic case) contributions, as happens at the two-derivative level. The covariant way of writing the Ricci scalar using the covariant derivative of $\tau_{\mu}{}^{a}$ and $\tau^{\mu}{}_{a}$ could be easily extended in order to incorporate the spin connection $\omega_{\mu ab}$, so that the action is also invariant under the $SO(1,1)$ transformations. In this way, the higher-derivative contributions could be written in a more compact form, since we expect large contributions if the curvatures are not chosen in an optimal way.
While DFT has been proved to be a powerful tool for reading the expansion of the A-field [22], taking the NR limit at the DFT level including higher-derivative terms is probably the best strategy to read these corrections and then parametrize them at the supergravity level. While in this paper we show that the Green-Schwarz mechanism for the gauge symmetry can be trivialized using field redefinitions, it is not clear that the same should happen for the gravitational one when addressing the $\alpha^{\prime}$-corrections for heterotic supergravity under the non-relativistic limit.
@@ Acknowledgements
The author is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
Appendix A Details on the cancellation of divergences when $\hat{A}_{\mu}$ is included
The field strength $\hat{F}_{\mu\nu}{}^{i}$, after expanding $\hat{A}_{\mu}{}^{i}$ considering (1.15), is given by
$\hat{F}^{(1)}_{\mu\nu}{}^{i}=\alpha_{-}^{i}\partial_{\mu}{\tau_{\nu}^{-}}+\partial_{\mu}{\alpha_{-}^{i}}\tau_{\nu}^{-}-\alpha_{-}^{i}\partial_{\nu}{\tau_{\mu}^{-}}-\partial_{\nu}{\alpha_{-}^{i}}\tau_{\mu}^{-}\,$, (A.1)
$\hat{F}^{(0)}_{\mu\nu}{}^{i}=+\alpha_{-}^{j}\eta^{ik}\tau_{\mu}^{-}a_{\nu}^{l}f_{jkl}-\alpha_{-}^{j}\eta^{ik}\tau_{\nu}^{-}a_{\mu}^{l}f_{jkl}\,$, (A.2)
$\hat{F}^{(-1)}_{\mu\nu}{}^{i}=\partial_{\mu}{a_{\nu}^{i}}-\partial_{\nu}{a_{\mu}^{i}}\,$, (A.3)
$\hat{F}^{(-2)}_{\mu\nu}{}^{i}=-\eta^{ij}a_{\mu}^{k}a_{\nu}^{l}f_{jkl}\,$, (A.4)
where the notation $\hat{F}^{(i)}_{\mu\nu}$ denotes the terms of order $c^{i}$ in the expansion. The contraction of this field strength with two inverse metrics, $\hat{F}^{\mu\nu i}=\hat{F}_{\rho\sigma}{}^{i}\hat{g}^{\mu\rho}\hat{g}^{\nu\sigma}$ yields the expression
$\hat{F}^{\mu\nu i}=\hat{F}^{(1)\mu\nu i}+\mathcal{O}\big(\frac{1}{c}\big)\,$, (A.5)
@@ Appendix A Details on the cancellation of divergences when $\hat{A}_{\mu}$ is included
$\hat{F}^{(1)\mu\nu i}=\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\sigma}h^{\nu\rho}\,.$ (A.6)
As the reader can easily check, the only contraction which produces non-vanishing divergences is given by $\hat{F}_{(1)\mu\nu}{}^{i}\hat{F}^{(1)\mu\nu j}\eta_{ij}$, which leads to the result (2.5).
Before analyzing the $\bar{H}$ squared contribution, we recall the form of the Chern-Simons terms,
$\bar{H}_{\mu\nu\rho}=3\left(\partial_{[\mu}\hat{B}_{\nu\rho]}-\hat{C}_{\mu\nu\rho}^{(g)}\right)\,$, (A.7)
$\hat{C}_{\mu\nu\rho}^{(g)}=\hat{A}^{i}_{[\mu}\partial_{\nu}\hat{A}_{\rho]i}-\frac{1}{3}\hat{f}_{ijk}\hat{A}_{\mu}^{i}\hat{A}_{\nu}^{j}\hat{A}_{\rho}^{k}\,.$ (A.8)
For the $\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}$ contributions, an important observation is that
@@
while the same object, when all the indices raised by the inverse metric, has the following form